OpenStudy (anonymous):

Can someone help me with this? 6logx-4logy-2logz write the expression as a single logarithm

Nnesha (nnesha):

r u familiar with log properties ??? @bluna6119

Nnesha (nnesha):

ello :-) ??

OpenStudy (anonymous):

not really.

Nnesha (nnesha):

quotient rule\[\huge\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\huge\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

first apply power rule and bec there are negative signs so use quotient rule